OpenStudy (anonymous):

HELP the equation for the directrix of this parabola is x=-1. x=1/4y^2 true or false

OpenStudy (anonymous):

not true

OpenStudy (anonymous):

\[\frac{1}{4}y^2= x\]\[y^2=1(4x)\]\[4p=1\]\[p=\frac{1}{4}\]focus = 1/4 dirextrix: x= -1/4 not x=-1

OpenStudy (anonymous):

I see 4p = 4 -> p = 1

OpenStudy (anonymous):

no

OpenStudy (anonymous):

if 4(x+1) than 4p =4 p=1 then directrix x=-1

OpenStudy (anonymous):

I'm confused, since I always remember y^2 = 4px

OpenStudy (anonymous):

yes but y^2 = 4x than 4x= 1(4x) 4p =1 p=1/4

OpenStudy (anonymous):

Thanks, Nancy! I'm looking it up now.

OpenStudy (anonymous):

what is the equation of the parabola, in vertex form, with focus at (2,-4) and directrix y=-6

OpenStudy (phi):

Click on properties in the box labeled "geometric figure"

OpenStudy (anonymous):

I check several other sources but it's not as the way the calculated, also the text book!!!

OpenStudy (anonymous):

help what is the equation of the parabola, in vertex form, with focus at (2,-4) and directrix y=-6

OpenStudy (anonymous):

Focus F (h, k+p) = (2, -4) Directrix y = k -p = -6 -> h = 2, k = 1, p = -5 => (x -2)^2 = 4(-5) ( y -1) Thus (x -2)^2 = -20 ( y -1)

OpenStudy (anonymous):

Rini, I'm going to fix the last one you asked me.